# Numerically “exact” charge transport dynamics in a dissipative electron–phonon model rationalizing the success of the transient localization scenario

### v1

@@ V Summary and outlook
In this study, we have considered an underexplored aspect of carrier transport in molecular semiconductors, namely, the influence of the damping of vibrational motions on transport dynamics. In contrast to many previous studies, which inferred transport properties from carrier dynamics starting from the factorized initial condition, our considerations are based on the numerically exact quantum dynamics of the current autocorrelation function of a one-dimensional dissipative model with nonlocal carrier–vibrational interaction. We use the DEOM representation of quantum dynamics, whose dissipaton algebra is essential to correctly treat the phonon-assisted current.
Whenever weakly damped (or undamped) vibrational dynamics are taken seriously, either numerically exactly or approximately, the diffusive carrier dynamics in the physically relevant high-temperature intermediate-interaction regime sets in from the superdiffusive side, giving rise to enhancements in the dynamical mobility below the vibrational frequency. Considering damping strengths characteristic of realistic materials, we find that the diffusive transport is approached from the subdiffusive side, and that the dynamical mobility exhibits no low-frequency upturns. Notably, this is in qualitative agreement with the majority of experimental results and the widely used transient localization scenario. Our findings explain the success of the exponentially decaying ansatz underlying the TLS in rationalizing experimental data. We suggest that low-frequency enhancements in the dynamical-mobility profiles in minimal models reflect the widely used assumption of undamped vibrations.
On the theoretical side, our study represents one of the pioneering applications of the DEOM formalism to computations of finite-temperature real-time correlation functions in dissipative electron–phonon models with more than few electronic levels and nonlocal electron–phonon interaction. We have also discussed the pitfalls of the widely used Drude–Lorentz bath model, which become apparent when considering mixed system–bath quantities, and how to avoid them. Thanks to the physically justified high-temperature assumption, our momentum-space formulation, and an appropriate DEOM closing with respect to the maximum depth $D$, we have succeeded in obtaining physically relevant results without resorting to hierarchy representations in nonstandard basis sets or using advanced DEOM propagation techniques. These could be exploited in further investigations of more realistic and challenging models, such as the multi-mode (instead of the single-mode) Brownian-oscillator model for the nonlocal interaction of the carrier with low-frequency intermolecular phonons or the explicit inclusion of the local interaction with high-frequency intramolecular vibrations.
@@ Acknowledgements.
$c_{m}=(-1)^{m+1}\frac{E_{0}\omega_{0}^{2}}{2\zeta}\left[\coth\left(\frac{i\beta\mu_{m}}{2}\right)-1\right]$ (50)
@@
In the Matsubara decomposition, $\xi_{m}=2\pi m,\rho_{m}=1$. In actual computations, we follow Ref. 101 and obtain $\xi_{m}$ and $\rho_{m}$ using the $(N_{\mathrm{BE}}-1)/N_{\mathrm{BE}}$ Padé approximant to the Bose–Einstein function. [102, 103]
@@ Appendix B Strong-damping limit
where $\Delta$ is given by Eq. (34) with $K=0$. Using Eq. (50) for $m=0$ and Eqs. (54) and (34), and keeping only the leading-order terms, we recast Eq. (55) as
@@
By representing Gaussian static disorder in terms of appropriate bosonic operators, Ref. 104 formally establishes that quantum dynamics of a carrier in the bath defined by Eq. (56) can be computed by averaging the dynamics modulated by the white noise in transfer integrals [the second term in Eq. (56)] over different realizations of Gaussian static disorder in transfer integrals [the first term in Eq. (56)]. $\mathcal{C}(t)$ in Eq. (11) and its approximate version Eq. (56) is the single-site bath correlation function, while nearest-neighbor transfer integrals are modulated by the difference between displacements of the corresponding local oscillators [Eq. (45)]. Therefore, the disorder strength is two times greater than in Eq. (56), i.e., $\sigma^{2}=2\times 2E_{0}T=2\lambda JT$ [Eq. (8)], which coincides with Eq. (46). The white-noise term ensures that the long-time diffusion constant remains nonzero for large but finite $\gamma_{0}$, which is nowadays known as the environment-assisted (or noise-assisted) quantum transport. [105, 106, 107] The TLS avoids computing (and then averaging) noise-modulated dynamics in many static-disorder realizations by resorting to Eq. (44), which approximates noise effects through the exponential correction to the dynamics in the static landscape.

### v2

@@ V Summary and outlook
In this study, we have considered an underexplored aspect of carrier transport in molecular semiconductors, namely, how a nonzero width of the spectrum of the vibrations modulating carrier’s hopping amplitude influences transport dynamics. Previous studies considered this aspect using the Drude–Lorentz spectral density and inferred transport properties from carrier dynamics starting from the factorized initial condition. Our considerations are instead based on the numerically “exact” quantum dynamics of the current autocorrelation function of a one-dimensional model with the physically more plausible (and formally more appropriate) Brownian-oscillator spectral density of nonlocal carrier–vibrational interaction. We use the DEOM representation of quantum dynamics, whose dissipaton algebra is essential to correctly treat the phonon-assisted current.
Whenever weakly damped (or undamped) vibrational dynamics are taken seriously, either numerically “exactly” or approximately, the diffusive carrier dynamics in the physically relevant high-temperature intermediate-interaction regime set in from the superdiffusive side, giving rise to enhancements in the dynamical mobility below the vibrational frequency. Considering damping strengths (widths of the relevant phonon spectrum) characteristic of realistic materials, we find that the diffusive transport is approached from the subdiffusive side and that the dynamical mobility exhibits no low-frequency upturns. Notably, this is in qualitative agreement with the majority of experimental results and the widely used transient localization scenario. Our findings explain the success of the exponentially decaying ansatz underlying the TLS in rationalizing experimental data. We suggest that low-frequency enhancements in the dynamical-mobility profiles in minimal models reflect the widely used assumption of a delta-like phonon spectrum (undamped vibrations).
On the theoretical side, our study represents one of the pioneering applications of the DEOM formalism to computations of finite-temperature real-time correlation functions in dissipative electron–phonon models with more than a few electronic levels and nonlocal electron–phonon interaction. We have also discussed the pitfalls of the widely used Drude–Lorentz bath model, which become apparent when considering mixed system–bath quantities, and how to avoid them. Thanks to the physically justified high-temperature assumption, our momentum-space formulation, and an appropriate DEOM closing with respect to the maximum depth $D$, we have succeeded in obtaining physically relevant results without resorting to hierarchy representations in nonstandard basis sets or using advanced DEOM propagation techniques. These could be exploited in further investigations of more realistic and challenging models, such as the multi-mode (instead of the single-mode) Brownian-oscillator model for the nonlocal interaction of the carrier with low-frequency intermolecular phonons or the explicit inclusion of the local interaction with high-frequency intramolecular vibrations.
@@ Acknowledgements.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@
$c_{m}=(-1)^{m+1}\frac{E_{0}\omega_{0}^{2}}{2\zeta}\left[\coth\left(\frac{i\beta\mu_{m}}{2}\right)-1\right],$ (50)
@@
In the Matsubara decomposition, $\xi_{m}=2\pi m,\rho_{m}=1$. In actual computations, we follow Ref. 102 and obtain $\xi_{m}$ and $\rho_{m}$ using the $(N_{\mathrm{BE}}-1)/N_{\mathrm{BE}}$ Padé approximant to the Bose–Einstein function. [103, 104]
@@ Appendix B Strong-damping limit
where $\Delta$ is given by Eq. (34) with $K=0$. Using Eq. (50) for $m=0$ and Eqs. (54) and (34) and keeping only the leading-order terms, we recast Eq. (55) as
@@
By representing Gaussian static disorder in terms of appropriate bosonic operators, Ref. 105 formally establishes that quantum dynamics of a carrier in the bath defined by Eq. (56) can be computed by averaging the dynamics modulated by the white noise in transfer integrals [the second term in Eq. (56)] over different realizations of Gaussian static disorder in transfer integrals [the first term in Eq. (56)]. $\mathcal{C}(t)$ in Eq. (11) and its approximate version Eq. (56) is the single-site bath correlation function, while nearest-neighbor transfer integrals are modulated by the difference between displacements of the corresponding local oscillators [Eq. (45)]. Therefore, the disorder strength is two times greater than in Eq. (56), i.e., $\sigma^{2}=2\times 2E_{0}T=2\lambda JT$ [Eq. (8)], which coincides with Eq. (46). The white-noise term ensures that the long-time diffusion constant remains nonzero for large but finite $\gamma_{0}$, which is nowadays known as the environment-assisted (or noise-assisted) quantum transport. [106, 107, 108] The TLS avoids computing (and then averaging) noise-modulated dynamics in many static-disorder realizations by resorting to Eq. (44), which approximates noise effects through the exponential correction to the dynamics in the static landscape.